What resolution is needed to separate ions with m/z values of 653.124 and 653.211?
Note: Use the mean m/z value for this calculation.

LOOKING FOR A NUMBER!! I know its a high resolution but give a specific answer!
Answer #1

Mean m/z = ( 653.124+653.211) /2 = 653.1675

delta m = difference between peaks = 653.211-653.124 = 0.087

Resolution = m /delta m = ( 653.1675) / ( 0.087) = 7507.6724